# Evidence-Decision-Feedback: Theory-Driven Adaptive Scaffolding for LLM Agents

### 24 Mar 2026 v2

@@ 5 Evaluation
To evaluate Copa, we investigate the following Research Questions:
@@
    Is students’ demonstrated understanding of the material aligned with their task mastery when interacting with Copa?
@@
    Is Copa’s feedback interpretable with respect to its evidentiary inferences?
In collaboration with five STEM educators and cognitive scientists (hereafter, “the educators”), we deployed Copa in an authentic high school science classroom in the southeastern United States with $n=33$ sophomore dyads [^2] We conducted all analyses at the dyad level (one student-pair per computer). (ages 15-16; 48% male/female, 2% no response; 68% White, 28% Hispanic, 20% Black, 12% Asian, 4% Native American) [^3] Several students identified as multiple ethnicities. participating in the C2STEM kinematics curriculum over six weeks. Sessions met once weekly for two hours and were taught by this paper’s lead author with oversight from the educators.
Students were first introduced to Copa five months prior to the study through a “teaser” activity that covered its pedagogical role, design rationale, and how it differs from tools such as ChatGPT; rather than providing direct solutions (ChatGPT), Copa supports students in constructing their own understanding. During the study, students completed two warm-up tasks to gain familiarity with Copa. The analyses in this paper focus on three subsequent tasks following this acclimation period. Each session began with instruction on 1- and 2-D kinematics, linking physics concepts (e.g., position, velocity, acceleration, $\Delta t$) with computational constructs (e.g., variable initialization and updates, loops, and conditionals), followed by computational modeling activities in C2STEM.
For this study, we collected 7,017 logged environment actions (mean $\approx 213$ per dyad per session), and 238 student-agent conversation turns (mean $\approx 7$ per dyad per session) across the three sessions, along with the CoT reasoning (i.e., “evidence” in Figure 3) associated with each agent decision. Together, these data streams constitute the primary dataset used for analysis in this paper, with all inferences manually triangulated using students’ conversations and screen recordings. Students completed an anonymous semi-structured exit survey. All findings were presented to three of the participating educators during a semi-structured Zoom discussion, providing additional feedback. Study participants provided informed consent/assent, student data were anonymized prior to analysis, and all procedures were approved by Vanderbilt University’s IRB.
@@ 5.1 RQ1: Scaffolding Adaptivity
As students achieve greater task mastery, Copa’s scaffolding is expected to adapt by gradually fading in accordance with ZPD principles. At the outset, Copa should prioritize probing students’ understanding to assess their current knowledge, approximate their ZPD, address prior misconceptions, and encourage critical thinking. As students deepen their understanding, Copa should look ahead, offering concrete suggestions to help them extend their reasoning beyond their current level of competence.
To evaluate Copa’s adaptivity (RQ1), we examined how its dialogue policies evolved within each session by analyzing the frequency with which policies were selected relative to quintiles of students’ task mastery percentages (see Table 1). We used quintiles for more stable comparisons of agent behavior across meaningful phases of learning. Policy frequencies were normalized as a percentage of the total dialogue within each dyad’s session, and then averaged across mastery quintiles. For each dialogue policy, we computed Spearman’s $\rho$ (due to the ordinal nature of the mastery data) to assess the strength of the relationship between shifts in dialogue policy and task mastery.
@@
As shown in Table 1, Copa significantly [^4] Throughout the paper, statistical significance is assessed at $\alpha=0.05$. reduced its use of the PROBE_UNDERSTANDING policy as mastery increases ($\rho=-0.34$, $p=0.034$), shifting toward implementation- and expansion-oriented support: SUGGEST_ACTION ($\rho=0.33$, $p=0.039$) and PUSH_LIMIT ($\rho=0.42$, $p=0.007$) increase in relative frequency. This demonstrates a pedagogical transition in which Copa moves from eliciting and assessing conceptual understanding to supporting the application of that understanding through model construction as mastery increases. This reflects ZPD alignment, as Copa shifts assistance toward challenges that extend learners beyond their independently-demonstrated competence, indicating Copa’s scaffolding appropriately adapted as task mastery increased.
@@ 5.2 RQ2: Understanding-Mastery Alignment
Students’ prioritization of task completion over conceptual understanding can present an overly optimistic picture of their learning, as students may appear to progress without being able to explain their code. From a social constructivist perspective, learning is evidenced not only by task performance but by learners’ ability to articulate and negotiate meaning through dialogue and explanation. Evaluating whether students’ task mastery is meaningfully connected to their verbalized understanding while interacting with Copa is, therefore, paramount.
To address RQ2, we examined the relationship between students’ response success rates—defined as the proportion of instances students correctly explained their code via a DEMONSTRATES_UNDERSTANDING dialogue state immediately following a PROBE_UNDERSTANDING dialogue policy—and students’ task mastery deciles. Deciles provide a finer-grained analysis of how students’ verbally demonstrated understanding varies across incremental levels of task mastery in C2STEM to assess alignment between the two. Data were normalized by averaging response success rate percentages for each dyad to mitigate interaction frequency bias. We again used Spearman’s $\rho$ due to the ordinal nature of the mastery deciles.
Figure 4: (a) Correlation between probing success rate and mastery deciles (RQ2).
@@
Figure 4a shows the success rate [^5] Success rate for mastery >0.7>0.7 was not calculated due to a lack of PROBE_UNDER-
STANDING policies: Copa chose to not probe knowledge once mastery was sufficient, as students stopped asking for direct answers and demonstrating misconceptions.
increases significantly as mastery increases ($\rho=0.40$, $p=0.014$), suggesting students are not merely “gaming the system”: mastery gains are accompanied by an improved ability to verbally reason about their code and respond correctly to Copa, i.e., C2STEM progress is meaningfully aligned with conceptual understanding rather than superficial task optimization.
Students’ environment actions advanced their model rather than deteriorated it, at approximately a 3:1 ratio, following agent interactions. The DEMONSTRATES_UNDERSTANDING dialogue state also showed a significant percentage increase as mastery improved ($\rho=0.34$, $p=0.030$), evidencing students’ increasing ability to articulate their reasoning as their problem-solving progressed. From a social constructivist perspective, articulating reasoning through dialogue reflects learners’ active construction of knowledge via interaction. Together, these findings demonstrate that students’ verbalization of their conceptual understanding aligned with their C2STEM task mastery.
@@ 5.3 RQ3: Student Reliance on Agent
Students’ control over how and when they interact with Copa risks overreliance. Ideally, as students approach task mastery, they would engage with Copa less frequently, reflecting increasing independence and reduced reliance on external support. Pursuant to SCT, this transition reflects the development of self-regulation and self-efficacy skills. However, prior work suggests this is not always the case, and students may develop a dependence on agent feedback that ultimately impedes their learning. This necessitates evaluating whether students use Copa in a manner consistent with productive support or continue to rely on the agent despite having developed sufficient task knowledge.
To evaluate the appropriateness of students’ reliance on Copa (RQ3), we examined how students’ voluntary engagement with the agent changed as they progressed through task mastery deciles. We computed the percentage of agent support (i.e., interactions) across mastery deciles to quantify the proportion of Copa turns per session. We again used deciles to provide fine-grained resolution over students’ progression toward mastery. To control for bias introduced by imbalanced interaction volumes, we normalized the data by averaging the agent-support proportion for each dyad across deciles, ensuring that high-frequency dyads did not exert disproportionate influence. We again used Spearman’s $\rho$ due to the ordinal nature of the mastery deciles.
Figure 4b shows students’ requests for support are highest early on, with approximately 59% of support occurring while mastery is low (<40<40%). As mastery increases, we observed a significant drop in requests for agent feedback (Spearman’s $\rho=-0.26$, $p<0.001$): students increasingly relied on their own knowledge, critical thinking, and collaboration with their partner rather than on agent support as mastery improved, consistent with developing learner autonomy, self-regulation, and self-efficacy—central SCT tenets.
@@ 5.4 RQ4: Feedback Interpretability
To trust agents, teachers, students, parents, and administrators must understand how they make decisions; otherwise, they risk limited adoption. EDF’s emphasis on interpretability aligns with ECD and Stealth Assessment, which require inferences about student knowledge be grounded in observable evidence collected unobtrusively during interaction. Agent feedback must be traceable to the dialogue policy that generated it, the evidence identified by the agent, and ultimately to the raw student data. To assess the interpretability of Copa’s feedback (RQ4), we decompose Copa’s internal reasoning chain (the agents’ “evidence” outputs in Figure 3) into the three links between EDF modules:
@@
Link 2, referred to as Alignment, assesses semantic coherence between Copa’s Decision (dialogue policy selection) and extracted Evidence. We compute sentence-level semantic similarity using Sentence-BERT embeddings (SBERT; all-MiniLM-L6-v2)—selected for its balance of efficiency and performance—by encoding Copa’s dialogue policy and dialogue state summaries, then comparing them via cosine similarity. Significance is evaluated via a permutation test (n=1000n=1000) [^6] Larger $n$ due to fewer utterances relative to environment actions in Link 1., comparing mean similarity of true evidence-policy pairs to shuffled baselines. Link 3, referred to as Faithfulness, evaluates whether Copa’s final Feedback adheres to its Decision. We assess faithfulness by again computing the SBERT-based cosine similarity between the dialogue policy and talk move, using the same permutation testing procedure as Link 2.
@@
Table 2: Interpretability results for all three links.
Table 2 shows that all three links exhibit statistically significant non-random structure ($p<0.001$), with similarity scores for true pairs approximately doubling those of the baselines. The Grounding score indicates Copa’s evidence extraction captured approximately twice as many meaningful tokens from the logs relative to the baseline (0.43 vs. 0.21). The Alignment score of 0.64 indicates strong semantic coherence between the extracted evidence and the dialogue policy selections, well above the baseline of 0.39. The Faithfulness score of 0.48 demonstrates that Copa’s generated responses remain semantically aligned with the dialogue policy, also outperforming the baseline (0.24). Viewed through the lens of ECD and Stealth Assessment, these results indicate Copa’s decisions are interpretable and grounded in explicit evidentiary reasoning.
@@ 5.5 Student and Teacher Perceptions
Post-study student survey responses revealed students responded positively to Copa overall, agreeing that “the questions Copa asked were appropriate” (mean =3.81=3.81 on a five-point Likert scale) and Copa “asks questions that made me think about my model” (mean =3.88=3.88). This suggests Copa successfully implemented its intended design goal of prioritizing critical thinking over direct solutions.
This behavior was not always appreciated. Ratings for perceived agent understanding (“I feel like Copa understood what I was doing”) and feedback utility (“I feel like Copa’s suggestions were useful”) were lower (mean =2.69=2.69 and $2.77$, respectively), indicating a disconnect between Copa’s pedagogical intent and students’ desire for answers. One group reported that Copa was “broken” and expected direct feedback rather than ZPD-aligned hints. This illustrates a broader tension between principled, scaffolded support and students’ expectations.
Findings aligned with educators’ classroom experiences. They noted that (1) students become frustrated when they do not receive immediate answers (“I agree with your conclusion…students wanting immediate answers”); (2) this expectation is shaped by their prior experiences with ChatGPT (“ChatGPT [is] so desperate to provide help…they’ve already become accustomed”); and (3) a persistent tension exists between instructional goals and student wishes, as students often “just [want]…the right answer…check the box and move on.”
One educator observed that student frustration was most pronounced due to repetitive interaction loops in which Copa withheld progress until understanding was demonstrated—particularly when students declined to act on feedback—noting that, while pedagogically intuitive, such questioning may not always be optimal [25]. Another viewed this positively, noting this behavior “gives Copa more of a human response”. Educators emphasized that students hold different expectations for agents than humans, underscoring the importance of student buy-in. To support this, one educator suggested sharing prior study findings with new cohorts at the outset to communicate Copa’s benefits and having Copa explain why students are stuck in such loops during interaction as a strategy for improving transparency and engagement.
@@ 6 Discussion and Conclusions
This work introduces an LLM-era, pedagogy-explicit scaffolding framework—Evidence-Decision-Feedback (EDF)—that treats adaptation and interpretability as first-class design goals. Operationalized through Copa, we demonstrated that an EDF-driven agent can adapt feedback to students’ evolving mastery in ways that are interpretable, avoid overreliance, and support the verbalization of conceptual understanding, consistent with social cognitive and social constructivist principles. Our modular, agentic design tightly couples learning theory with system architecture and is evaluated in an authentic classroom setting, rather than solely through benchmarks or single-shot feedback.
Theory emphasizes that effective support should help students build understanding rather than just providing answers. However, consistent with prior work [7], our findings indicate that students do not always value this approach. Research also emphasizes the importance of learner agency for engagement [5], raising questions about the appropriate degree of autonomy when it conflicts with established pedagogical principles. Authors et al. (2025) [7] argue that student agency should be supported only to the point that it does not undermine critical thinking opportunities, a view supported by our findings. Overall, our results indicate that while EDF allows Copa to apply learning theories effectively, they reveal a tension between principled guidance and students’ preferences.
From a methodological perspective, our results emphasize adaptation as a first-class system capability, requiring continuous scaffold adjustments based on learner evidence rather than executing fixed strategies. Recent work on agentic AI failures found that agents often appear effective during testing but degrade in real use due to a lack of adaptation [16]. This reinforces the need to evaluate agentic systems in authentic classroom settings, where learners’ evolving behaviors, needs, and goals reveal limitations not visible in controlled benchmarks.
Practically, effective deployment of agentic systems in situ requires attention beyond system design. As Thomas et al. (2025) [29] note, the value of agent feedback ultimately depends on students’ willingness to engage with it. Our findings reinforce the need for the AIED community to invest in fostering student buy-in and trust, as meaningful learning benefits are unlikely if students disengage from or resist agent support. Deeper engagement is, therefore, a prerequisite for realizing the pedagogical potential of LLM agents in practice.
Our limitations include evaluating scaffold adaptation as a function of task mastery rather than a broader range of attributes, and the absence of an RCT due to institutional constraints. Reported relationships are correlational, not causal, and may be influenced by students’ prior knowledge and interaction dynamics. However, correlational evidence has been shown to be valuable for informing student performance, understanding, and the role of pedagogical agents in authentic settings [14]. We note (for reference only) that students using Copa exhibited higher normalized learning gains (0.53) than the prior cohort not using an agent (0.45) from the same school, curriculum, and demographics.
Overall, our work demonstrates that (1) agentic, multi-agent LLM systems can be designed to operate in principled, interpretable, and pedagogically meaningful ways; and (2) EDF offers a viable framework for operationalizing learning theory within LLM-based agents towards these ends. In future work, we will expand our adaptive scaffolding evaluation beyond task mastery to include factors like prior knowledge, self-regulation, self-efficacy, and social dynamics. We hope this work inspires further AIED research that moves beyond surface-level LLM interactions toward theoretically grounded, adaptive, and trustworthy educational agents that can meaningfully support learning at scale.

### 29 Mar 2026 v3

@@ 5 Evaluation
To evaluate Copa, we address the following Research Questions:
@@
    Is students’ verbalized understanding aligned with their task mastery when interacting with Copa?
@@
    Is Copa’s feedback interpretable with respect to its evidentiary inferences and student input data?
In collaboration with five STEM educators and cognitive scientists (hereafter, “the educators”), we deployed Copa in a high school science classroom in Nashville (USA) with $n=33$ sophomore dyads [^2] We conducted all analyses at the dyad level (one student-pair per computer). (ages 15–16; 48% male/female, 2% no response; 68% White, 28% Hispanic, 20% Black, 12% Asian, 4% Native American) [^3] Several students identified as multiple ethnicities. participating in the C2STEM kinematics curriculum over six weeks. Sessions met once weekly for two hours and began with instruction on 1- and 2-D kinematics—linking physics concepts (e.g., position, velocity, acceleration, $\Delta t$) with computational constructs (e.g., variable initialization and updating, loops, and conditionals)—followed by computational modeling tasks in C2STEM. Classes were taught by this paper’s lead author with oversight from the educators. Our analyses focus on three tasks: a 1-D truck, a 2-D drone dropping one package, and a 2-D drone dropping two packages to different targets.
We collected 7,017 logged environment actions (mean $\approx 213$ per dyad), and 238 student-agent conversation turns (mean $\approx 7$ per dyad) across the three sessions, along with the CoT reasoning (i.e., “evidence” in Figure 3) associated with each agent decision. Students completed an anonymous exit survey. All findings were presented to three of the participating educators during a semi-structured Zoom discussion to solicit additional feedback. Together, these data constitute the primary dataset for our analyses, with all inferences manually triangulated through student conversations and screen recordings. Study participants provided informed consent/assent, student data were anonymized prior to analysis, and all procedures were approved by Vanderbilt University’s IRB.
@@ 5.1 RQ1: Scaffolding Adaptivity
As students achieve greater task mastery, Copa’s scaffolding must adapt by gradually fading in accordance with ZPD principles. At the outset, Copa should prioritize probing students’ understanding to assess their current knowledge, approximate their ZPD, and address prior misconceptions. As students deepen their understanding, Copa should look ahead and offer concrete suggestions to help them extend their reasoning beyond their current level of competence.
To evaluate Copa’s adaptivity (RQ1), we examined how its dialogue policies evolved within each session by analyzing the frequency with which policies were selected relative to quintiles of students’ task mastery percentages (see Table 2). We used quintiles for more stable comparisons of agent behavior across meaningful phases of learning. Policy frequencies were normalized as a percentage of the total dialogue in each dyad’s session and then averaged across mastery quintiles. For each dialogue policy, we computed Spearman’s $\rho$ (due to the ordinal nature of the mastery data) to assess the strength of the relationship between shifts in dialogue policy and task mastery.
@@
As shown in Table 2, Copa significantly ($\alpha=0.05$) reduced its use of the PROBE_UNDERSTANDING policy as mastery increased ($\rho=-0.34$), shifting toward implementation- and expansion-oriented support: SUGGEST_ACTION ($\rho=0.33$) and PUSH_LIMIT ($\rho=0.42$) increased in relative frequency. This demonstrates a pedagogical transition in which Copa moves from eliciting and assessing conceptual understanding to supporting its application through model construction as mastery increases—reflecting ZPD alignment as Copa shifts assistance toward challenges beyond learners’ current competence and indicating Copa’s scaffolding appropriately adapted as task mastery increased.
@@ 5.2 RQ2: Understanding-Mastery Alignment
Students’ prioritization of task completion over conceptual understanding can present an overly optimistic picture of their learning, as students may appear to progress without being able to explain their code. From a social constructivist perspective, learning is evidenced not only by task performance but by learners’ ability to articulate and negotiate meaning through dialogue and explanation. Evaluating whether students’ task mastery is meaningfully related to their verbalized understanding while interacting with Copa is, therefore, paramount.
Figure 4: (a) Correlation between success rate44footnotemark: 4 and mastery deciles (RQ2).
@@
footnotetext: Copa chose not to probe students’ knowledge if mastery was greater than 0.7.
To address RQ2, we examined the relationship between students’ response success rates—defined as the proportion of instances students correctly explained their code via a DEMONSTRATES_UNDERSTANDING dialogue state immediately following a PROBE_UNDERSTANDING dialogue policy—and students’ task mastery deciles. Deciles provide a finer-grained analysis of how students’ verbally-demonstrated understanding varies across incremental levels of task mastery to assess alignment between the two. Data were normalized by averaging response success rate percentages for each dyad to mitigate interaction frequency bias. We again used Spearman’s $\rho$ due to the ordinality of deciles.
Figure 4a shows the success rate increased significantly as mastery increased ($\rho=0.40$), suggesting students were not merely “gaming the system”: mastery gains were accompanied by an improved ability to verbally reason about their code and respond correctly to Copa, i.e., C2STEM progress was meaningfully aligned with conceptual understanding rather than superficial task optimization. Pursuant to Social Constructivism, articulating reasoning through dialogue reflects learners’ active knowledge construction through interaction. Together, these findings demonstrate that students’ verbalized conceptual understanding aligned with their C2STEM task mastery.
@@ 5.3 RQ3: Student Reliance on Agent
Students’ control over when and how they interact with Copa promotes agency but risks overreliance. This can impede persistent learning and necessitates evaluating whether students use Copa in a manner consistent with productive support or continue to rely on the agent despite having developed sufficient mastery. As students progress, they should engage Copa less often, reflecting increasing independence and reduced reliance on external support. Pursuant to SCT, this transition reflects the development of self-regulation and self-efficacy skills.
To evaluate the appropriateness of students’ reliance on Copa (RQ3), we examined how students’ engagement with the agent changed across task mastery deciles. We computed the percentage of agent support (i.e., interactions) across deciles to quantify the proportion of Copa turns per session. To control for bias introduced by imbalanced interaction volumes, we normalized the data by averaging the agent-support proportion for each dyad across deciles. We again used Spearman’s $\rho$ due to the ordinal nature of deciles.
Figure 4b shows students’ requests for support were highest early on, with approximately 59% of support occurring while mastery was low (<40<40%). As mastery increased, we observed a significant drop in requests for agent feedback (Spearman’s $\rho=-0.26$): students increasingly relied on their own knowledge and collaboration with their partner rather than on agent support as mastery improved, consistent with developing learner autonomy, self-regulation, and self-efficacy—central SCT tenets.
@@ 5.4 RQ4: Feedback Interpretability
Agents must be transparent about how they make decisions in order for stakeholders to trust them; otherwise, they risk limited adoption. EDF’s emphasis on interpretability aligns with ECD and Stealth Assessment, which require inferences about student knowledge to be grounded in observable evidence collected unobtrusively during interactions. Agent feedback must be traceable to the dialogue policy that generated it, the evidence identified by the agent, and ultimately to the raw student input data. To assess the interpretability of Copa’s feedback (RQ4), we decompose Copa’s internal reasoning chain (the agents’ “evidence” outputs in Figure 3) into the three links between EDF modules:
@@
Link 2, referred to as Alignment, assesses semantic coherence between Copa’s Decision (dialogue policy selection) and extracted Evidence. We compute semantic similarity using Sentence-BERT embeddings (SBERT; all-MiniLM-L6-v2)—selected for its balance of size and performance—by encoding Copa’s dialogue policy and dialogue state summaries, then comparing them via cosine similarity. Significance is evaluated via a permutation test (n=1000n=1000) [^5] Larger $n$ due to fewer utterances relative to environment actions in Link 1., comparing mean similarity of true evidence-policy pairs to shuffled baselines. Link 3, referred to as Faithfulness, evaluates whether Copa’s final Feedback adheres to its Decision. We assess faithfulness by again computing the SBERT-based cosine similarity between the dialogue policy and talk move, using the same permutation testing procedure as Link 2.
@@
Table 3: Interpretability results for all three EDF links.
Table 3 shows that all three links exhibit statistically significant non-random structure ($p<0.001$), with similarity scores for true pairs approximately doubling those of the baselines. The Grounding score indicates Copa’s evidence extraction captured approximately twice as many meaningful tokens from the logs relative to the baseline (0.43 vs. 0.21). The Alignment score of 0.64 (vs. 0.39) indicates strong semantic coherence between the extracted evidence and the dialogue policy selections. The Faithfulness score of 0.48 (vs. 0.24) demonstrates that Copa’s generated responses remain semantically aligned with the dialogue policy. Viewed through the lens of ECD and Stealth Assessment, these results indicate Copa’s decisions are interpretable with respect to how observable learner data is transformed into evidentiary representations, pedagogical decisions, and student-facing feedback.
@@ 5.5 Student and Teacher Perceptions
Post-study survey data revealed students responded positively to Copa overall, agreeing that “the questions Copa asked were appropriate” (mean =3.81=3.81 on a five-point Likert scale) and Copa “asks questions that made me think about my model” (mean =3.88=3.88). This behavior was not always appreciated, as ratings for perceived agent understanding (“I feel like Copa understood what I was doing”) and feedback utility (“I feel like Copa’s suggestions were useful”) were lower (mean =2.69=2.69 and $2.77$, respectively). While students reported appreciating Copa’s inquiry-based approach, they often became frustrated with the agent when their desire for direct answers was not met. This highlights a broader tension between principled support and students’ agency and expectations.
Our findings aligned with educators’ experiences, who agreed that (1) students become frustrated when they do not receive immediate answers (“I agree with your conclusion…students wanting immediate answers”)—an expectation shaped by students’ prior ChatGPT experience, they noted; and (2) a persistent tension exists between instructional goals and student wishes, as students often “just [want]…the right answer…check the box and move on.” They also remarked that students hold different expectations for agents than humans, underscoring the importance of students “buying in” to agent instructional intent.
@@ 6 Discussion and Conclusions
In this paper, we presented EDF as a theory-driven adaptive scaffolding framework for LLM agents. Through Copa, we demonstrated EDF enables agent feedback that reflects students’ evolving mastery while remaining interpretable, avoiding overreliance, and supporting knowledge verbalization in line with social-cognitive and social-constructivist principles. Copa’s modular, agentic design provides a tight theory-to-architecture coupling, which we evaluated in an authentic classroom setting. Overall, we showed that (1) agentic, multi-agent LLM systems can be designed to operate in principled, interpretable, and pedagogically meaningful ways; and (2) EDF offers a viable framework for operationalizing learning theory within LLM-based agents toward these goals.
Theory emphasizes that effective support should help students build understanding rather than just providing answers. However, consistent with our prior work, students do not always value this [9]. Their frustration when denied direct answers reveals a design tension we term the helpfulness paradox: pedagogical agents must be helpful enough to earn trust and invite use when needed, yet restrained enough to avoid encouraging cognitive offloading that harms durable learning. Effective LLM agents must therefore balance immediate utility with constraints that preserve independent problem solving.
From a methodological perspective, LLM pedagogical agent adaptation requires continuous scaffold adjustment based on learner evidence rather than fixed strategy execution. Recent work on agentic AI failures shows that agents often appear effective in testing but degrade in authentic use when they fail to adapt [18]. This further motivates evaluation in real classroom settings, where learners’ shifting behaviors, needs, and goals expose limitations that controlled benchmarks often miss. Practically, however, real-world impact depends not only on system design but also on students’ willingness to engage with agent support. As Thomas et al. (2025) [34] note, the value of agent feedback is ultimately constrained by whether students choose to use it. Fostering student buy-in and trust is therefore essential, as learning benefits are unlikely to emerge if students disengage from or resist agent feedback.
Our findings also point to a broader implication for evaluation: agents must be assessed across multiple dimensions. Prior agent evaluations often center on a single learner or agent attribute—e.g., theoretical alignment, learning gains, self-efficacy [10]—rather than evaluating agents holistically. In our study, adaptivity, understanding-mastery alignment, reliance, interpretability, and stakeholder perceptions are complementary, each capturing important, interconnected aspects. Accordingly, multi-lens agent evaluation must become the norm rather than the exception, as performance gains along a single dimension may obscure failures in other dimensions that are equally critical to meaningful learning.
The primary limitations of our work include evaluating scaffold adaptation solely as a function of task mastery rather than a broader set of learner attributes, and the absence of an RCT due to institutional constraints: reported relationships are correlational rather than causal. Our evaluation focuses on within-system behavior rather than comparison against alternatives, and although EDF is a general framework, we evaluate it through only one implementation, Copa. Researcher-led classroom instruction may introduce effects that would not arise under independent teacher implementation. Manual triangulation across log, screen, and video data suggests Copa supports self-regulation and produces pedagogically sound reasoning chains; however, rigorous stakeholder-centric evaluation across these axes remains future work.
This paper’s central contribution is demonstrating that EDF can be used to operationalize LLM agents in pedagogically meaningful ways, as supported by our findings. In future work, we will address the limitations above by broadening evaluation to include self-regulation, self-efficacy, social dynamics, and in-depth stakeholder interpretations. Our findings and future research emphasize the importance of grounding LLM agents in learning theory, empirical evidence, stakeholder perspectives, and robust evaluation. We hope this work motivates further AIED research that moves beyond surface-level LLM interactions toward more adaptive and trustworthy agents that can support learning at scale.